1.      How many ml of 1.00M HCl do I need to add to 10.0ml of 1.00M NaOH to get a solution with pH = 13?


1
Expert's answer
2021-10-17T11:21:04-0400

Let the volume of HCl added = VmL

Moles of NaOH = (10mL)(1.00M) = 10 mol

Moles of HCl = (VmL)(1.00M) = V mol

pH = 13 or pOH = 1 (solution is basic or excess of NaOH is left)

or, [OH-] = 10–1 = 0.1

Excess NaOH left = (10-V) mol

Final volume of the solution = (V+10)mL

Final [OH-] = (10-V)/(V+10) = 0.1 (calculated above)

or, 10-V = 0.1V + 1

9 = 1.1V

V = 8.18 mL